OpenStudy (idily101):

what is 4% of 30 and show your work

OpenStudy (anonymous):

1.2, put x over 30=4 over 100. cross multiply for x. you get 120=100x. solve for x.... check my work?

OpenStudy (anonymous):

what is = x is= equals 4%=4% of= multiply 30=30 x=.04 x 30 x=1.2

OpenStudy (anonymous):

4% of 30= 4/100*30 =.04*30=1.2 .